OpenStudy (anonymous):

Given: tan theta=-5/12 and theta is in Quadrant IV. cos theta= ?

OpenStudy (anonymous):

\[\cos \theta = 12/13\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

welcome :)